40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 19 23:36:19 2012, yuta, Update, LSC, calibrated POY error signal LSCyarmTF_usingITMY.pngLSCyarm_HAover1plusG.pngPOYerrorcalibration.png
    Reply  Wed Jun 20 00:01:04 2012, Jenne, Update, LSC, calibrated POY error signal LSC_POY_11_I_ERR_calib_19June2012.pdfPOY_calib_19June2012_FiltBankSettings.png
    Reply  Wed Jun 20 18:43:57 2012, yuta, Update, LSC, calibrated POX error signal LSCxarmTF_usingITMX.pngLSCxarm_HAover1plusG.pngPOXerrorcalibration.png
Message ID: 6835     Entry time: Wed Jun 20 00:01:04 2012     In reply to: 6834
Author: Jenne 
Type: Update 
Category: LSC 
Subject: calibrated POY error signal 

[Yuta, Jenne]

We have measured the out of loop residual motion of the Yarm while locked with the ALS.  We see ~70pm RMS, as compared to Kiwamu's best of ~24pm RMS.  So we're not yet meeting Kiwamu's best measurement, but we're certainly not in crazy-land.

The Yarm ALS was locked, I took a spectrum of POY11_I_ERR, and used the calibration that we determined earlier this evening.  For reference, I attach a screenshot of our ALS loop filters - we had on all the boosts, and both resonant gain filters (~3Hz and ~16Hz).

A large part of the RMS is coming from the 60Hz power line and the 180Hz harmonic....if we could get rid of these (how were they eliminated from the measurement that Kiwamu used in the paper?? - plotted elog 6780) we would be closer. 

Also, it looks like the hump (in our measurementf ~100Hz, in Kiwamu's ~200Hz) is not quite an order of magnitude higher in amplitude in our measurement vs. Kiwamu's.  We have ~5e-11 m/rtHz, Kiwamu had ~7e-12 m/rtHz.  This increase in noise could be coming from the fact that Yuta and Koji decreased the gain in the Ygreen PDH loop to prevent the PDH box from oscillating. 

While we should still think about why we can't use the same gain that Kiwamu was able to ~6 months ago, we think that we're good enough that we can move on to doing mode scans and residual motion measurements of the Xarm.

 

Attachment 1: LSC_POY_11_I_ERR_calib_19June2012.pdf  14 kB  | Hide | Hide all
LSC_POY_11_I_ERR_calib_19June2012.pdf
Attachment 2: POY_calib_19June2012_FiltBankSettings.png  34 kB  | Hide | Hide all
POY_calib_19June2012_FiltBankSettings.png
ELOG V3.1.3-